OpenStudy (osuxbucks15):

Viola drives 200 meters up a hill that makes an angle of 9° with the horizontal. To the nearest tenth of a meter, what horizontal distance has she covered? 1,262.8 m 200.4 m 197.5 m 31.3 m

OpenStudy (anonymous):

 \(\cos(9)=\frac{x}{200}\)

OpenStudy (anonymous):

and so \(x=200\cos(9)\) and a calculator

OpenStudy (anonymous):

...a calculator in degree mode to be precise ;-)

OpenStudy (osuxbucks15):

? i dnt get it

OpenStudy (anonymous):

yes, degree mode for sure